OpenStudy (anonymous):

Find the measures of two angles, one positive and one negative, that are coterminal with pi/5

OpenStudy (anonymous):

coterminal... does that mean that the two angles add to pi/5?

OpenStudy (anonymous):

i honestly dont know

OpenStudy (anonymous):

I just looked it up. http://hotmath.com/hotmath_help/topics/coterminal-angles.html

OpenStudy (tkhunny):

Just add and subtract \(2\pi\) \(\pi/5 + 2\pi\;and\;\pi/5 - 2\pi\) A little algebra to pretty it up and you'll be done.

OpenStudy (anonymous):

and now that you've had the answer dumped on you...

OpenStudy (anonymous):

do you understand why?

OpenStudy (anonymous):

Honestly kinda. But i would like to fully get it

OpenStudy (tkhunny):

No dumping, just a target to shoot for. Plenty of algebra practice to go.